OpenStudy (anonymous):

Shana wants to buy a new bicycle that has a retail price of $259.99. She knows that it will be on sale next week for 30% off the retail price. If the tax rate is 7%, find the total amount, to the nearest cent, that she will save by waiting until next week.

OpenStudy (anonymous):

To get the discount amount, you multiply the retail price by 0.3 since that's the discount. To get the tax amount, multiply the retail price by 0.07. f(x) = 259.99 - 0.3x + 0.07x f(259.99) = 259.99 - 0.3(259.99) + 0.07(259.99)

OpenStudy (anonymous):

Can you take it from there?

OpenStudy (anonymous):

i dont understand because i am taking a practice regents and it says the correct answer is $83.46. how did they come up with this answer?

OpenStudy (anonymous):

It says what she will SAVE.

OpenStudy (anonymous):

son how do you get that?

OpenStudy (anonymous):

so**

OpenStudy (anonymous):

83.46? I don't see how. Hold on for a second.

OpenStudy (anonymous):

yes here is the link http://www.nysedregents.org/IntegratedAlgebra/112/ialg12012-rg.pdf its number 39

OpenStudy (anonymous):

Alright. I see what happened.

OpenStudy (anonymous):

First, apply the 7% tax to the 259.99 and add it to it.

OpenStudy (anonymous):

259.99(1.07) = 278.1893 Then, you have to find the discount on the retail price. 259.99 - 259.99(0.3) = 181.997 Then apply the tax to this new price. 181.997(1.07) = 194.73251 Then find the difference to get the amount saved. 278.1893 - 194.73251 = 83.45679 which rounds to $83.46

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

yes i understand it thank you so much for all your help but just one quick question how does the tax become 1.07 from .07?

OpenStudy (anonymous):

Basically, by multiplying by 0.07 gives you the tax amount and then you'd add it to the original. However, my multiplying by 1.07, it adds it automatically. :) This is only if you have to add the extra amount.

OpenStudy (anonymous):

Less plugging into a calculator.

OpenStudy (anonymous):

ohh ok. thank you so much for all your help tonight i appreciate it very much :)

OpenStudy (anonymous):

You're welcome :)